# Think or Step-by-Step? UnZIPping the Black Box in Zero-Shot Prompts

### v2

@@ A.4 Human Evaluation
A.5 Qualitative Analysis
Task Information MSW LLM Answer with Original Prompt (Partial) Replaced Word LLM Answer with Perturbed Prompt (Partial)
True Answer: 2412
Prompt: 0-IRR
Dataset: GSM8k
Irrelevant 1. Each member pays an annual fee of \$150 for snacks.
2. Each member buys 6 hardcover books at \$30 each: 6×30 = 180
3. Each member buys 6 paperback books at \$12 each: 6×12 = 72
Next, we find the total amount collected from all 6 members: 6×402 = 2412
The final answer is 2412.
Unrelated 1. Each member pays:
- Annual fee for snacks: \$150
- Cost for six hardcover books: \$30
- Cost for six paperback books: \$12
Total per member: 150 + 30 + 12 = 192
2. There are 6 members, so the total amount collected by Niles is: 192×6 = 1152
The final answer is 1152.
Alice, Bob, Claire, Dave, and Eve are playing a game. …. Finally, Dave and Alice swap balls. At the end of the game, Bob has the yellow ball. Options: (A) yellow ball. (B) white ball. (C) red ball. (D) pink ball. (E) blue ball.
True Answer: A
Prompt: 0-CoTR
Dataset: Big Bench
Right Bob has the yellow ball.
Correct answer is (A).
- Bob has the yellow ball.
Correct answer is (B).
True Answer: 336
Prompt: 0-CoTB
Dataset: AQUA-RAT
step-by-step 1. Calculate the number of apples and bananas in 4 dozen:
- 1 dozen = 12
- Therefore, 4 dozen = 4 × 12 = 48 pieces of fruit.
2. Determine the cost of apples and bananas:
- The cost of one banana is Rs. 3.
- The cost of one apple is Rs. 4.
3. Calculate the total cost for 48 fruits:
- Cost of 24 bananas = 24 × Rs. 3 = Rs. 72.
- Cost of 24 apples = 24 × Rs. 4 = Rs. 96.
4. Add the costs together:
- Total cost = Cost of bananas + Cost of apples
- Total cost = Rs. 72 + Rs. 96 = Rs. 168.
- 1. Cost of One Banana: Rs. 3
2. Cost of One Apple: Rs. 4
Next, we calculate the cost for 4 dozen (which is 48) of each fruit:
- Cost of 4 dozen Bananas:
Cost of 4 dozen bananas = 48 bananas × 3 Rs/banana = 144 Rs
- Cost of 4 dozen Apples:
Cost of 4 dozen apples = 48 apples × 4 Rs/apple = 192 Rs
Total cost = Cost of bananas + Cost of apples = 144 Rs + 192 Rs = 336 Rs
Table 16: Comparison of GPT-4o mini’s responses to original and perturbed prompts in classification tasks. Each row shows how changes to the most significant word (MSW) through synonym/co-hyponym replacement or removal affect model output.
Original Sentence True Translation Prompt Translation Language MSW LLM translation with original prompt Replaced word LLM translation with perturbed prompt
Der Vater unseres Landes wusste etwas von schlechter Presse. The father of our country knew something about bad press. 0-DTG German to English Refine The father of our nation knew something about bad press. Enhance The founding father of our nation was aware of the negative publicity.
 {CJK*}UTF8gbsn推动工作的新思路新举措新成效 To promote new ideas, new actions, and new effects of the work. 0-DPS Chinese to English Translation promote new ideas, new measures, and new results in work. Paraphrase advancing efforts with innovative concepts, new initiatives, and achieving new outcomes.
Table 17: Comparative analysis of original and perturbed prompt responses by GPT-4o mini on selected translation task instances where modifications to the most significant word (MSW) of the instructional prompt are made through synonym or co-homonym replacement.
Prompt: 0-IRR Dataset: GSM8k [Uncaptioned image]
Prompt: 0-CoTR Dataset: Big Bench [Uncaptioned image]
Prompt: 0-CoTB Dataset: AQUA-RAT [Uncaptioned image]
Table 18: The ZIP score heatmaps of the instructional prompts applied to the classification datasets featured in the qualitative analysis.
Prompt: 0-DTG Language: German to English [Uncaptioned image]
Prompt: 0-DPS Language: Chinese to English [Uncaptioned image]
Table 19: The ZIP score heatmaps of the instructional prompts applied to the translation datasets featured in the qualitative analysis.
@@ A.7 Perturbations
Let’s work this out in a
step-by-step way
to be sure we have the
right answer.
Let’s work this out in a
step-by-step way
to be secured we have the
right answer.
Let’s work this out in a
step-by-step way
to be we have the
right answer. (removal)
Let’s work this out in a
step-by-step way
to be concrete we have the
right answer.
Let’s work this out in a
step-by-step way
to be certain we have the
right answer.
Let’s work this out in a
step-by-step way
to be secured we have the
right answer.
Let’s work this out in a
step-by-step way
to be we have the
right answer.
Let’s work this out in a
step-by-step way
to be concrete we have the
right answer.
Let’s work this out in a
step-by-step way
to be certain we have the
right answer.
Let’s work this out in a
step-by-step way
to be we have the
right answer.
Let’s work this out in a
step-by-step way
to be certain we have the
right answer.
Let’s work this out in a
step-by-step way
to be certain we have the
right answer.
@@
Let’s first understand the
problem and devise a
plan to solve the problem.
Then let’s carry out the
plan and solve the problem
step-by-step.
Let’s first understand the
problem and devise a plan
to tackle the problem.
Then let’s carry out the plan
and solve the problem
step-by-step.
Let’s first understand the
problem and devise a plan
to cope the problem.
Then let’s carry out the plan
and solve the problem
step-by-step.
Let’s first understand the
problem and devise a plan
to crack the problem.
Then let’s carry out the plan
and solve the problem
step-by-step.
Let’s first understand the
problem and devise a plan
to tackle the problem.
Then let’s carry out the plan
and solve the problem
step-by-step.
Let’s first understand the
problem and devise a plan
to cope the problem.
Then let’s carry out the plan
and solve the problem
step-by-step.
Let’s first understand the
problem and devise a plan
to crack the problem.
Then let’s carry out the plan
and solve the problem
step-by-step.
Let’s first understand the
problem and devise a plan
to tackle the problem.
Then let’s carry out the plan
and solve the problem
step-by-step.
Let’s first understand the
problem and devise a plan
to cope the problem.
Then let’s carry out the plan
and solve the problem
step-by-step.
Let’s first understand the
problem and devise a plan
to tackle the problem.
Then let’s carry out the plan
and solve the problem
step-by-step.
Table 22: Example of perturbation filtering for classification prompts. Bold words perturbed via synonyms, co-hyponyms, removal. Filtered by semantic similarity (>30%) and grammar. Rejections in red.
Zero-shot
Instructional Prompt
Generated Candidates Semantic Similarity >30%
Meaningful and
grammatically correct
Final Perturbations
Please provide the translation
for the following sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the upcoming sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the after sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the next sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the consecutive sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the upcoming sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the after sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the next sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the consecutive sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the consecutive sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the consecutive sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the next sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the upcoming sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the next sentence
step-by-step and then provide
the complete sentence.
@@
Table 23: Example of perturbation filteringg process for translation task prompts. Bold words are perturbed via synonyms, co-hyponyms, and removal. Perturbations are filtered by semantic similarity (>30%) and grammatical correctness. Rejections in red.
@@
Table 24: Evaluation of semantic similarity thresholds (20%-50%) on the 0-PS prompt using 20 manually validated word variants. Each row shows an original word, its generated alternatives with similarity scores, and acceptance decisions across thresholds. Green indicates correct decisions (accept/reject), red indicates incorrect ones. The accuracy row demonstrates that 30% provides optimal performance (95%) in identifying meaningfully related perturbations.
@@ A.8 Token Usage per Instructional Prompt
Table 25: Estimation of token usage per instructional prompt across different datasets (AQUA-RAT, Big Bench, and GSM8K) using GPT-4o mini. This table presents the input tokens, output tokens, and total tokens calculated with space-based tokenization. Actual tokenization methods employed by GPT-4o mini may vary, and thus these figures serve as approximations of token usage.

### v3

@@ A.5 Human Evaluation
0-CoT 0-CoTB 0-CoTR 0-IRR 0-PS
Participant 1
Think
step-by-step
Work
Problem
step-by-step
Work
Ignore
Irrelevant
Information
Understand
Problem
Plan
Participant 2
step-by-step
step-by-step
step-by-step
Right
Answer
Ignore
Irrelevant
Devise
Plan
step-by-step
Participant 3
Think
Work
Right
Irrelevant
Solve
Participant 4
step-by-step
step-by-step
step-by-step
Irrelevant
step-by-step
Participant 5
step-by-step
Breath
step-by-step
step-by-step
Sure
Right
Ignore
Irrelevant
Description
Understand
Plan
Solve
Table 17: Important words identified by participants across five zero-shot prompts for classification tasks.
@@ A.7 Perturbations
Table 20: Illustration of the multi-stage filtering of prompt perturbations for classification prompts. Each candidate is generated via synonym, co-hyponym, or removal, and filtered for semantic similarity (>>30%) and grammaticality. Underlined candidates indicate perturbations that were rejected at a given filtering stage. The final column shows the valid perturbations used for evaluation.
@@
Table 21: Evaluation of semantic similarity thresholds (20%–50%) on the 0-PS prompt using 20 manually validated word variants. Each row lists an original word, its generated candidate replacements (with similarity scores), and whether the candidate is accepted or rejected at each threshold. Bold and underlined entries indicate incorrect acceptance/rejection (contradict human judgment). The accuracy row shows the overall agreement rate with human judgments for each threshold. The 30% threshold yielded the best performance (95% accuracy), which is why it was chosen for the main experiments.
@@ A.8 Token Usage per Instructional Prompt
Table 22: Estimated token usage per instructional prompt across datasets using GPT-4o mini. Input, output, and total tokens are approximated using space-based tokenization and may differ from GPT-4o mini’s actual method.